What numbers add to get -14 and multiply to get 49

Answer by stanbon(75887)

What numbers add to get -14 and multiply to get 49
----
Equations:
x + y = -14
xy = 49
----------------
Rearrange:
x = -14-y
-----
Substitute:
(-14-y)y = 49
-14y - y^2 = 49
-----
y^2 + 14y + 49 = 0
Factor:
(y+7)^2 = 0
---
y = -7 with multiplicity 2.
----------------------------
Answer: -7 and -7
=====================
Cheers,
Stan H.